OpenStudy (anonymous):

What is the equation of the line that passes through the point (2, 3) and is perpendicular to the line y = –x – 3? y = –x – 1 y = –x + 5 y = x + 1 y = x + 5

OpenStudy (anonymous):

step 1) when two lines are perpendicular, the product of their slopes is "-1" use this information and find the slope of the new line m = ?

OpenStudy (anonymous):

@some_someone

OpenStudy (anonymous):

@electrokid im not sure what m equals

OpenStudy (anonymous):

what is the slope of the given line \(y=-x-3\)

OpenStudy (anonymous):

1st, find the slope, m1, of y = –x – 3 by putting it in slope-intercept form. (That means solve for y) In this case its already in slope intercept form. Slope intercept form: y = mx + b

OpenStudy (anonymous):

m1 = -1 --------- The slope of lines perpendicular is the negative inverse of m1, = +1

OpenStudy (anonymous):

Use y = mx + b and the point to find b. 3 = (1)(2) + b 3 = 2 + b 1 = b

OpenStudy (anonymous):

y = x + 1 is thru (2 , 3) and perpendicular to y = –x – 3

OpenStudy (anonymous):

so @some_someone is the answer c?

OpenStudy (anonymous):

Precisely :D

OpenStudy (anonymous):

thank you can you help me with the next answer please